X^2 - y^2 = cx


1
Expert's answer
2022-06-06T16:07:25-0400

Differentiate both sides with respect to "x"

"2x-2y\dfrac{dy}{dx}=c"

Then


"2x^2-2xy\dfrac{dy}{dx}=cx"

Substitute


"2x^2-2xy\dfrac{dy}{dx}=x^2-y^2"


"x^2+y^2-2xy\dfrac{dy}{dx}=0"